OpenStudy (anonymous):

A cube has a surface area of 600 square meters. What is its volume? 100 ft3 300 ft3 600 ft3 1000 ft3

Parth (parthkohli):

\[600 = 6s^2\]Solve for \(s\), then find \(s^3\)

OpenStudy (anonymous):

okay ...give me a sec

OpenStudy (anonymous):

600=6s^2 600/6=s^2 100=s^2 \[\sqrt{100}=\sqrt{s^2}\] \[\sqrt{100}=10 s=10

Parth (parthkohli):

Yes!

Parth (parthkohli):

Now because the volume is \(s^3\), you need to find \(10^3\).

OpenStudy (anonymous):

1000

Parth (parthkohli):

That's the volume :-)

OpenStudy (anonymous):

Thank you!!